OpenStudy (meilendurcer):

Please I really need help with this... Award will be given. 2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 over 8 = 7x−1over 6 24 over 1(3x+6 over 8) = 24 over 1(7x−1 over 6) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

\[\frac{ 3x+6 }{ 8 } = \frac{ 7x-1 }{ 6 } \frac{ 24 }{ 1 } (\frac{ 3x+6 }{ 8 } ) = \frac{ 24 }{ 1 } (\frac{ 7x-1 }{ 6 } )\]

OpenStudy (meilendurcer):

yeah that

OpenStudy (anonymous):

is it two equations or just one? \[\frac{ 3x+6 }{ 8 } = \frac{ 7x-1 }{ 6 } \] '[\frac{ 3x+6 }{ 8 } = \frac{ 7x-1 }{ 6 } \] \[\frac{ 24 }{ 1 } (\frac{ 3x+6 }{ 8 } ) = \frac{ 24 }{ 1 } (\frac{ 7x-1 }{ 6 } )\] [\frac{ 24 }{ 1 } (\frac{ 3x+6 }{ 8 } ) = \frac{ 24 }{ 1 } (\frac{ 7x-1 }{ 6 } )\]

OpenStudy (anonymous):

oh ok, do you see in the third step on the right side when we have 4(7x-1) [if we distribute the 4, we would have...] 28x-4 how is this different then what is shown in the problem?

OpenStudy (meilendurcer):

in the problem its 28x-1

OpenStudy (anonymous):

right, do you see how Max forgot to do the 4*1 part and just carried the 1 down from the line above?

OpenStudy (meilendurcer):

yes

OpenStudy (anonymous):

so, we have found the error Max made. and now you can 'explain it to him' the only part left of the problem is to tell him how he can check his answer if he did it correctly the first time. Max got "x=1" as his answer. the method to check if it is correct is to plug in 1 for every x he see's in the original equation and see if the equation is still true.

OpenStudy (meilendurcer):

ok thanks a lot!

OpenStudy (meilendurcer):

what about the left side. same thing?

OpenStudy (anonymous):

the left side is all correct, there aren't any mistakes made. can you finish this question and do that part, "provide the correct solution"

OpenStudy (meilendurcer):

with some help i think

OpenStudy (anonymous):

so we got this far: 3(3x+6) = 4(7x−1) 9x+18 = 28x−4 the next step is to move all the x-terms to the same side and move all the number terms to the opposite side.

OpenStudy (meilendurcer):

ok i'm with you o far

OpenStudy (anonymous):

can you try doing the next step? we aren't allowed to do your homework, but we can help

OpenStudy (meilendurcer):

do I substitute the -9x?

OpenStudy (anonymous):

you will want to move the 9x to the other side by subtracting both sides by -9x, good. so what does that look like now?

OpenStudy (anonymous):

@DemolisionWolf Did you ever get a awnser back?